Santa Clara University Robotic Systems Laboratory
GeneSat1 Operations
 
DISCLAIMER
This is NOT a NASA site.  We intend to provide information that is accurate; however, we do not guarantee this accuracy.  The contents of this site are being provided as a way to share the excitement and progress of this mission with the public.
 
CURRENT MISSION STATUS

Mission Phase = Post-Experiment Phase

Major Events to Date

12/16/06 0400 PST Minotaur successfully launched from WFF 
12/16/06 0420 PST Deployment of GeneSat from PPOD
12/16/06 0530 PST Beacon signal received at SRI station and by amateurs
12/17/06 0400 PST 2.4 GHz command channel functional
12/18/06 0500 PST GeneSat biological experiment initiated
12/22/06 0535 PST 96 hour biological experiment complete and all baseline data downloaded.
12/22/06 0635 PST Data distributed to mission team scientists and engineers within 60 min of experiment conclusion (and within 30 min of download), and plots of primary science data publicly distributed within 90 min of experiment conclusion (and within 60 min of download).
1/17/07 1700 PST All operations for executing the primary mission criteria have been successfully performed, with results disseminated.
2/21/07 1430 PST Handover of mission responsibility from NASA Ames to Santa Clara University.  Initiation of the mission's post-experiment phase.
SATELLITE STATUS
In general, the satellite has maintained a nominal health condition for all parameters (to include temperature, power, CPU performance, etc.) over the duration of the primary mission.  The 2.4 GHz command link is functional but limited in performance; the amateur radio beacon has experienced some drift, but is functional.  Temperature control on the payload is degrading, but this function is no longer in use for mission purposes.
  
GROUND SEGMENT STATUS 
All Control Nodes have nominal functionality - SRI is currently the active Control Node.
SRI Communication station is nominal although it is reconfigured intermittently for other experiments or for routine maintenance.
SCU Communication station is nominal.
 
BEACON STATUS - NOMINAL
 The beacon is generally functional although there has been some drift of the main frequency.  The beacon is occasionally off in order to support other operations.  Beacon Users - e-mail the ops team if you are interested in using the beacon and would like to know the operating schedule.

 

Mission Operations Log Summaries

Jan 08 - Contacts run on 1/8, 1/9, and 1/16.  Beacon on, bus health nominal.  S-Band comm signal strength data being collected.  Payload heater set-point changed to 16 C (from 20C) and no difference in payload temp control behavior was observed (50 mA current draw remains rather than an expected value on the order of 500-700 mA).  Long term studies show little to no performance loss of solar panel power over the past year.  Still no CPU resets.  Historical bus and payload data downloaded as time permits (several hundred pages have been downloaded over the past few months).

Nov-Dec 07 - Contacts run on November 14, 16, 27, and 29 as well as on Dec 11,14, and 17.  Satellite bus health continues to be nominal, and the payload temperature control anomaly continues (this anomaly has no effect on operations at this point in the mission). The 12/14 contact was unsuccessful due to comm station equipment having been disconnected by non-SCU personnel in order to support other operations at the SRI station; there was not time to reconfigure the station properly prior to the planned contact.  Giovanni is our current operations lead, and he is collecting long-term platform performance data that is supporting the evolution of the bus design for future missions (we have 3 planned at this point in 2008) as well as the development of a future publication.  Note - 12/16/07 was the 1-year birthday of GeneSat in orbit!

10/16/07 - The ops team was finally able to get back into the SRI station although some work remains to be done (like re-installing the UHF feed). Two contacts were run today, lde by the fearless I. Mas.  The beacon was re-enabled and verified, and state of health data was downloaded given the extended time over the past 2 months during which no operations were run.  Satellite bus health is nominal, and the previously observed payload temperature control anomaly continues (again, we note that the anomaly is for a function that is no longer needed to support the mission).  A new ops trainees participated in the contacts and showed off by hand-tracking the beacon with a hand-held antenna and transceiver (nice job Jose!).  Post-pass logging was initiated via the internet from Munich! We currently plan to run the next set of contacts next week, and the ops team is working with the engineering team to establish a set of end-of-life test objectives. CK

Mid-August through Mid-October Summary - The SRI station is unavailable during this time for maintenance, retrofit, and other test operations.  The beacon will be shut off during this period. CK

Early August 2007 Ops Summary - The ops team is training several new mission controllers for educational purposes as well as initial training for the team's next satellite, PharmaSat.  The team is retrieving routine health data and is also conducting power stress tests (which occasionally require the beacon to be turned off) in order to assess the power generation capabilities of the system.  The satellite bus health is nominal. Payload health is good with the exception of the payload temperature control process (temp is below setpoint, the control loop is activated, but a much lower than nominal heater current draw exists); we will monitor this system and develop a plan to evaluate the heater FET and/or other possible root causes.  No operations during the week of Aug 12 given the attendance of most of the team at the AIAA Conference on Small Satellites (which includes presentation of several GeneSat-1 related papers). CK

July 2007 Ops Summary - Only limited operations conducted in July due to routine maintenance with the SRI station and its use to support other studies.  Satellite beacon off. CK

May 2007 Ops Summary - The team continues to support MAST, although contact has been lost and support for that mission may be ending soon.  Health and training contacts continue.  In addition, initial power stress testing initiated in order to provide platform performance data to support design work on the team's next spacecraft, PharmaSat (which uses an extended version of the GeneSat-1 bus).  The beacon service is intermittent due to this testing. CK

April 2007 Ops Summary - Limited operations conducted given the ops team's support of the Tethers Unlimited Inc MAST tethered satellite mission.  MAST uses the SRI station as well as SCU's 2.4 GHz radio equipment.  See the MAST site for details.  GeneSat-1 remains in nominal condition, and the limited contacts that are being performed include standard health analyses as well as practice operations for a student laboratory class.  Satellite beacon shut off near end of the month. CK

3/29/07, 0830 PST -  The beacon has been re-enabled and the vehicle is in nominal SOH.  For the time being, communication link tests have been completed (and the results will be published in a paper to be published later this year). Satellite health remains nominal. CK

2/25/07 - 3/25/07 -  NOTE - the ops team is conducting a series of communication link tests during this next month in order to better characterize the operation of the command transceiver.  During this time, the beacon may be turned off.  CK

2/21/07, 1430 PST -  We are now in the mission's Post-Experiment Phase!  In today's ceremony at 1430 PST, mission responsibility for GeneSat-1 was handed over from NASA Ames to Santa Clara University.  Santa Clara will continue to operate the mission until the satellite re-enters or becomes non-functional.  Ops activities will continue with a focus on bus characterization, support of student research and educational experiments, use of the vehicle to support crew training exercises for the next mission (PharmaSat), support of education and public outreach activities, etc. CK

2/20/07, 1200 PST -  Satellite health nominal. Less than ideal, but still functional, communications on both the 2.4 GHz link as well as the beacon link continue with the cause believed to be related to the vehicle's attitude profile. CK

2/13/07, 2300 PST -  During this evening's contact at ~2200 PST, control node operations were run from the Ames MMOC.  Vehicle health was verified as nominal: no cpu resets, internal temperature at 18.85 degrees C, external temperatures entering eclipse of 11.8-13.1 degrees C, mean payload temperature at 25.27 degrees C (set-point at 25 degrees C), and payload pressure at 14.32 psi.  The beacon signal was strong.  Command channel success rate was 27%.  Archived biology experiment data was downloaded.  CK

2/7-12/07 -  SRI station pointing recalibration is being performed during this period.  The next scheduled pass will be in the evening of 2/13/07. CK

2/6/07, 0920 PST -  During the second contact, transceiver link testing was conducted (so, no standard commands sent or telemetry received).  The link was closed with pointing offsets on the order of 0.46 deg AZ and 0.72 deg EL once pointing hit 16 EL ascent. Margins were in the 20 dB to 0 dB range with a period of 80-90 sec and a 40 sec period modulation on top of that (as previously observed).  The beacon was audibly faint throughout the contact, but only a few packets were successfully decoded.  Troubleshooting will continue. CK

2/6/07, 0740 PST -  During today's first contact (~0630 PST), the 2.4 GHz link was closed with a 32% command channel success rate.  Vehicle health was verified as nominal (in eclipse: solar panel temps -1-0 deg C, internal temp 13.4 deg C, no cpu resents, payload temp 25.36 deg C for 25 deg C setpoint, 14.126 psi). Only a few beacon packets were received. CK

2/5/07, 1530 PST -  No contacts were performed today.  The team installed a new 2.4 GHz transceiver at the SRI station to address any possible issues with that component.  In addition, SRI personnel started a new dish calibration process to verify pointing accuracy. CK

2/2/07, 1000 PST -  Three contacts were executed (~0530 - 0900 PST) in order to more fully characterize the sporadic communications issues that have been occurring. Prior to the first contact, a full reboot of the ground transceiver was performed - this solved the intermittent link issues with the ground loop-back test.

During the first contact, only 1 audible beacon packet was received - at 7 EL ascent (the team later ran an orbit analysis and this seemed to correspond to the time of eclipse exit).  No 2.4 GHz commands were successfully executed.  

During the second contact, no beacon was received, and no 2.4 GHz commands were successfully executed.

During the third contact, 2.4 comm signal strength tests were performed.  The link was closed for a portion of the contact (closed at 21 EL on ascent through max EL of 24 deg, then a single test packet at 18 EL on descent), but the signal strength was about 10 dB below the (nominal) values collected during the previous week.  No beacon at either the SRI or SCU station until the very end of the contact at about 5 deg EL descent - 3 packets heard at SRI and 1 at SCU. 

The team will run multiple contacts again on Tuesday; in addition, several checkouts of the station will be performed to address any potential ground segment issues contributing to this issue.  We note, however, that we are seeing anomalous behavior now on two comm channels, and the beacon issues are being observed through two stations. CK

2/1/07, 1130 PST -  Three SCU beacon station checks were conducted today with no beacon reception at all. CK 

1/31/07, 1000 PST -  Once contact run at ~0850 PST.  Good beacon throughout contact, but no commands successfully executed through the 2.4 GHz link; the lack of 2.4 GHz success is possibly related to a ground transceiver issue in which problems with the ground loopback testing has been noticed; the ops team will be back on Friday to look into this issue further. CK

1/30/07, 1130 PST -  Two  contacts were executed today at ~0805 PST and ~0945 PST. During the first contact, beacon reception was limited again (AOS at 30 EL ascent and lost at 15 EL descent; generally weak signal; no reception through SCU station, suggesting a beacon issue vs a ground segment issue for a single point anomaly assumption).  Standard command and telemetry operations were conducted, with ~ 60% success rate and the command link closed from 45 EL ascent through 10 EL descent). Vehicle health verified as nominal (battery voltage 8.51, external solar panel temperatures from 0-2 deg C, internal temp 12.5 deg C, no cpu resents, payload temp 24.79 deg C for 25 deg C setpoint, pressure 14.13 psi).  

During the 2nd contact, no beacon was received (at either station).  The first part of this contact was run with standard command/telemetry operations in order to collect health data that could illuminate the beacon issues.  During the second portion of the pass, the transceiver was switched to test mode to continue with 2.4 GHz transceiver/link performance testing - great data received on that - plots show the spacecraft's coning motion!  CK

1/29/07, 1000 PST -  One contact was performed in order to perform link testing - good signal strength data was collected during the contact.  The beacon was strong throughout and showed good vehicle health (payload temp at 25.27 degrees C given 25 deg C setpoint, internal satellite temp at 15 deg C, microhard current being drawn nominally, etc.). CK

1/26/07, 1300 PST -  One contact was run today at ~1225 PST in order to perform radio link testing (meaning no standard command and telemetry operations).  Unfortunately, the team was unable to close the link with the transceiver in test mode; the team will be re-evaluating the test configuration.  Also, the beacon signal was noticed to be weak, with acquisition not occurring until ~ 29 EL on ascent; beacon continued through max EL but then was lost at ~ 38 EL on descent.  Potential linkage between the two comm issues, but not enough time to evaluate; the team will continue to track this issue on the next contact. CK

1/25/07, 1400 PST -  One contact was run today at ~1330 PST.  Vehicle health was verified as nominal (battery voltage 8.11, external solar panel temperatures 8-12 degrees C, internal temperature 17.6 degrees C, no cpu resets, payload temp controlled at 25.58 degrees C (setpoint at 25), and pressure 14.34 psi), and additional 96-hour biological experiment data was retrieved.  Command channel success rate was just shy of 70%.  Minor SRI station maintenance is being performed on a daily basis with no expected interference with realtime operations.  The team is preparing to run a series of link analysis tests over the next few weeks; this will be performed by putting the 2.4 GHz transceiver into a test mode, thereby allowing signal strength to be logged during a contact. CK

1/23/07, 1330 PST -  One contact was run today at ~1005 PST.  Vehicle health was verified as nominal, and additional 96-hour biological experiment data was retrieved.  The ops team is preparing for a transfer of the control node to the Ames Bldg 240 Ops Center, with a target date of Feb 1.  CK

1/18/07, 1630 PST -  One contact was run today at ~1555 PST.  Vehicle health was verified as nominal, and additional 96-hour biological experiment data was retrieved.  CK

1/17/07, 1700 PST -  Beacon is back on.  The ops crew completed the 3rd contact of the day (and the 4th in the overall procedure for collecting high-resolution solar panel current data).  Execution of the procedure was a success (even Dave, the project's systems engineer seems pleased!).  Vehicle health remains nominal.  The team has now successfully completed all operations required to declare that the the primary mission is a success (of course, we have many secondary objectives that will now be performed over the rest of the mission).  CK

1/17/07, 1400 PST -  We have just run the first of several contacts today as part of a continuing procedure to collect high-resolution solar panel current data.  As part of this procedure, the beacon has just been turned off; it will be re-enabled later today. Commanding success took a nosedive compared to yesterday - the success rate was ~ 43% and commanding wasn't really reliable until about 40 degrees EL.  CK

1/16/07, 1630 PST -  The ops team is back in action after a wonderfully restive long weekend.  One contact was run this afternoon.  Vehicle health was verified as nominal, and additional experiment data was retrieved.  We also set the payload temperature control set-point to 15 deg C (it had been at 25 deg C); this was done as the first step in a multi-contact, two-day procedure to capture another collection of high-resolution solar panel current data.  The remainder of this procedure will be executed during several sequential contacts tomorrow afternoon.  HAMs be warned - the beacon will be off for a portion of this time.  Great news - the command channel success rate jumped to 79.6%, our highest ever (and a shot of redemption for command operator Mike).  As a test for the next period of operations, the command and telemetry operators conducted this contact were running ops out of the Primary MOC in Ames Bldg 240.  An additional operator was operating out of the SCU GeneSat lab, and of course, the tracking operator was at SRI.  All functionality good running in full distributed operations mode (although we need to upgrade that speakerphone in 240!).  Facilities note - the Ames Primary MOC was down from 0500-0600 PST this morning due a scheduled Ames firewall reboot; there was no affect to realtime operations.  CK

1/11/07, 1730 PST -  All is well.  We ran one contact today at ~ 1650 PST.  Vehicle health is nominal across the board.  We set the payload temperature control set-point to 25 degrees C.  We also continued to improve the resolution of data for the 96 hour experiment.  The ops team will stand down until next Tuesday.  On Tuesday or Wednesday (Jan 16 or 17), we are currently planning on running another multi-contact procedure to download more high-resolution solar panel current data.  When we do this, we will be temporarily disabling the beacon again - we will post details as we near that time.  We've had more than 23,000 beacon packets received by the amateur radio community!  CK
Ladies... would you date this man?  Neither would we... but he did get certified to run the dish!

1/10/07, 2030 PST -  Attention to HAMs - you may notice odd beacon data at this point.  Telemetry originating from the bus should be nominal, but those points originating from the payload may bounce all over the place.  This is expected; it most likely started around noon PST today.  Here's the convoluted explanation.... Our memory for storing experimental payload data has run out.  We will ultimately be able to overwrite this memory with new experiment data (meaning optics data since the biology is all dead at this point), but we need to clear a flag to allow this to happen.  We won't send that command until we get everything the everybody/anybody wants from that bank of memory (this includes the critical 96 hour experiment data plus lots of data after that).  While we're in this state, we are not updating payload data (primarily since we not collecting most of that data for storage); the effect is that this information is undefined - and it appears that the undefined state is noise, or gibberish, or whatever you want to call it.  Anyway - thanks to those of you who've contacted us with questions and heads-up's about this issue - we hope this explains what's going on.  We'll let you know when we clear the flag.  In other news, we ran 2 contacts this evening - bus side is great.  Also - congratulations to Paul and Giovanni for getting formally certified to run tracks on the SRI station!  CK

1/9/07, 2012 PST -  Beacon is back on!  CK

1/9/07, 1915 PST -  We are in the process of running a multi-contact procedure to collect a special set of solar panel current data.  This will be high-resolution data (1/sec) over a period of about 17 min, scheduled to be performed in full sunlight.  Our hope is that this extended sampling will give us an improved sense of our attitude profile, g-level, etc.  This test is one of the last major data collections required for declaring a success for the primary portion of the GeneSat-1 mission.  FYI - the reason that the beacon was disabled during this process was due to wanting to have as predictable and continuous a current draw as possible.  So, the beacon and microhard components were turned off.  In addition, the payload temperature control system was commanded to attempt to draw a maximum amount of current during this time.  To do this, the control setpoint was set to 15 deg C last night, and then the set-point was raised to 34 deg C immediately prior to the   
sampling; the idea is that this will cause the control heaters to be on at max power during the majority of the sampling.  The beacon and microhard will be re-enabled once the collection is over.  During our last contact at approx 1830 PST, we did much of the commanding to set up the data collection.  During the next contact at approx 2010 PST, we will reconfigure comm and start downloading data.  So far, so good.... CK.

1/9/07, 1820 PST -  NOTE TO HAMS - we are about to disable to beacon for an orbit in order to do a special data collection.

1/8/07, 1830 PST -  One contact was performed at ~ 1740 PST in order to a) perform a vehicle state of health, b) complete the download of high resolution payload data from the end of the 96-hr experiment when the temperature control set-point was changed from 34 degrees C to 25 degrees C, c) set the payload temp to 15 degrees C in anticipation of tomorrow’s special solar panel current data collection, and d) continue improving the 96 hour payload data resolution with an “every 2 pages” collection policy.  The contact was a success.  

Satellite health is nominal.  Payload temperature control was holding at the set-point of 25 degrees C, but was reset to a set-point of 15 degrees C.  This commanded set-point change was functionally verified through the set-point telemetry point in the payload register file.  By the end of the contact, the mean payload temperature was observed to be ~22 degrees C; the ops team will continue to track this change tomorrow.  Payload pressure is 14.358 and humidity is 92.94%.  Bus health is excellent with steady temperatures (interior temp on external payload canister is 17.5 degrees C, solar panel external temperatures are in the 17.5 to 21.8 degrees C range early in eclipse), bus power is good (battery voltage is at ~9.8V and all power lines are at their expected levels), and there have been no CPU resets. 

Command channel communications remain functional with a command channel success rate of 50.8%.  We note that tonight’s ops were one of the first time where we had reasonably good 2.4 comm throughout the entire duration of the antenna trajectory (given that it is only 50% successful).  Coincidently, this occurred on the same day that NORAD officially converted our TLE designation from “OBJECT C” to “GENESAT” – we’re now official!  Beacon downlink remains strong with packets audible and decoded throughout the entire contact (even before track is started and after it is terminated at 5 degrees EL). 

Ground segment status is nominal.  Upgrade and repairs at the SRI station have been completed; in addition to addressing all wind damage issues, the SCU team also installed a new light on the tripod (which greatly assists in visually monitoring the antenna through a remote camera feed).  There were no antenna glitches tonight!  The ops team continues to practice distributed operations as a team to include data exchange via the ops line, etc.  

In amateur radio news, more than 21,400 beacon packets have been submitted by external operators.  Also, SCU students ran a successful test this evening, attempting to capture beacon packets with a handheld radio and 5-element hand-pointed antenna (using a very ad-hoc pointing technique).  7 packets were received!  This was about 30 sec worth of data – this result is quite promising, and given this success we will explore a better pointing process for successive contacts.  CK.

1/6/07, 2000 PST -  One contact was run this evening in order to perform a vehicle state of and to download higher resolution payload data from the end of the 96-hr experiment when the temperature control set-point was changed from 34 degrees C to 25 degrees C.  The contact was a success, although another contact will need to be run next week to complete the download of this high resolution data.  Satellite health is nominal.  Payload temperature control is holding at the current set-point of 25 degrees C (mean temp is 25.38 degrees C).  Payload pressure is 14.358 and humidity is 92.938%.  Bus health is excellent with steady temperatures (interior temp on external payload canister is 17.99 degrees C, solar panel external temperatures are in the 17 to 19 degrees C range early in eclipse), bus power is good (battery voltage is at ~8.3V and all power lines are at their expected levels), and there have been no CPU resets.  Command channel communications remain functional although still below the desired level.  Our command channel success rate for this evening was 25%.  Beacon downlink remains strong with packets audible and decoded throughout the entire contact (even before track is started and after it is terminated at 5 degrees EL).  We will not run contacts tomorrow.  Next week, we will be attempting to complete the download of mission data required to declare a mission success.  This data will include the high-resolution temperature transient data at the end of the 96 hour experiment as well as a new non-standard collection of high resolution solar panel current data (to assist in our assessment of the spin rate, g-levels being experienced by the payload, etc.).  No antenna glitch tonight! CK

1/5/07, 1200 PST -  No operations today due to scheduled maintenance on dish (assessment and workarounds for the servo glitching problems we've been having) and the desire for a crew down day due to holiday travel, etc.  CK

1/4/07, 1800 PST -  This evening's contact was cancelled due to high winds. CK

1/3/07, 1900 PST -  One contact was run this evening in order to verify satellite health and to improve the data resolution of the early portion of the biological experiment.  Satellite health remains nominal.  7 commands were successfully processed at a command channel success rate of 25% when the AZ antenna servo shut down at about zenith due to a peak gust of wind (so we had to terminate the track).  The good news, however, is that Paul and Giovanni repaired the mesh damage this afternoon in record time (so, of course, you know the quality of work must have been outstanding.... probably as good as the original installation job..... which couldn't take a mild hurricane force wind.....).  Anyway, the original forecast was that the job might take several days, so we're using the opportunity to check out and possibly upgrade/tweak a few other parts of the station.  Right now, Mike Cousins is out in the utility shack checking breakers to see if we can curtail the servo cut-off glitches that we've been seeing.  CK

1/2/07, 2145 PST -  Happy New Year! Two contacts were run today to verify health, to retrieve high resolution solar panel current, and to continue downloading low-resolution post-experiment payload data. 

The first contact had a very high EL (84.4 deg) so track was going to have to be terminated half way through (do to AZ rates that exceed the capabilities of the station); unfortunately, an odd timing problem occurred with the dish's programmed track control software (one which we had not personally seen before but which has been reported to occasionally happen).  We had to terminate track, but Giovanni used the blown contact to practice tracking the satellite manually - wasn't good enough to attempt commanding, but it was an interesting practice session and he was able to catch the beacon on the descent.  After the contact, we troubleshot the problem by simulating the pass but we couldn't reproduce the problem; GPS timing, pointing angles, etc all made sense.  Oh well... 

So, we ran a second contact 90 min later. That was a success.  Satellite health is in outstanding shape - nominal across the board as reflected by the dashboard. The command channel success rate was 37%. 

Dish repairs are scheduled to start tomorrow and, weather permitting, will continue for the rest of the work week.  The dynamic duo of master antenna refurbishers (Paul and Giovanni) will reprise their roles in this encore performance. CK

12/31/06, 0245 PST -  Two contacts were executed today at ~ 2335 and 0110 PST.  During these contacts, the ops team ran standard state of health procedures, and then used their remaining time to gather low-resolution post-experiment payload health, to gather additional high-resolution payload health data from the first few hours of the experiment, and to obtain high-resolution bus data for the most recent orbit. Satellite health remains nominal.  Payload temperature is holding steady at its set point of 25 degrees C, with sensor values in the 23.61 to 26.67 deg C range.  Payload pressure is 14.383, and humidity is 88.175%.  Bus power is nominal, there have been no CPU resets, and temperatures are as expected (the external payload enclosure temp is ~ 19.4 deg C, and solar panel temps early in eclipse are ~ 10 to 11 deg C). Ground segment remains functional and nominal with the exception of the SRI antenna mesh damage caused by the severe windstorms on 12/26-27/06.  The command channel success rate for today’s 2 contacts was approximately 44%.  The NORAD-supplied TLEs continue to change noticeably with each daily update, and those used for this day’s contacts provided much better performance than the first set from Day 14. The ops team continues to practice telephone protocols with distributed team members in preparation for a smooth transition of the control node to non-SRI locations.  More than 17,200 beacon packets have been submitted to date by external amateur radio operators.  The team will not run contacts on Jan 1 due to a scheduled holiday break for New Year’s. CK

12/30/06, 0100 PST -  We ran 2 contacts this evening and were able to retrieve a reasonable amount of low-resolution post-experiment payload data from the past week as well as high-resolution bus health data over the previous orbit.  Satellite health remains nominal.  Payload temperature is holding steady at its set point of 25 degrees C, with sensor values in the 23.04 to 25.98 deg C range.  Payload pressure is 14.466, and humidity is 91.545%.  Bus power is nominal, there have been no CPU resets, and temperatures are as expected (the external payload enclosure temp is ~ 18 deg C, and solar panel temps early in eclipse are ~ 15 to 16 deg C).  Our command channel success rate continues to be low (34%), almost certainly due to the damaged mesh.  However, we note that the TLE's seemed noticeably poor on our first contact; coincidently, NORAD issued a new set in time for our second contact, and performance significantly improved.  We plan on running a contact tomorrow evening, and then taking 2 days off for the New Year's holiday. CK

12/28/06, 2330 PST -  Believe it or not, even with some significant damage to the dish's mesh, we were able to get a few commands through to the vehicle in this evening's first contact.  Our second contact was terminated early due to an AZ servo shutdown - a common occurrence, although this was the first time it caused us to terminate an active track (other than during the windstorm).  The few commands we were able to send showed that the satellite remains healthy with the payload temperature being successfully controlled at 25 degrees C.  During this evening's contacts, the ops team also practiced telecon protocols in preparation for running contacts with the control node at a location other than at the SRI communication station. CK

12/28/06, 1530 PST -  High winds yesterday damaged the mesh on the dish (as well as a lot of other facilities/items in the foothills such as the mesh on the 150 ft dish, power lines, etc.); SFO reported gusts up to 65 mph, and the foothills often sees gusts at a little above the SFO level. We ran inspections and functional checkouts today to confirm that everything other than the physical mesh seems to be OK. Given the holidays, restoration of the mesh will probably have to wait until next week.  We will be trying a contact later today to see if 2.4 GHz communication is even possible with the current damage. CK

Wind damage to the dish - but it is still good enough to support both 2.4 GHz as well as beacon communications!

12/27/06, 0930 PST -  High winds (with gusts at 50+ mph) have motivated us to cancel any contacts for today.  The wind is affecting antenna operation.  Given that the primary experiment data has already been retrieved, we've decided to cancel today's contacts and resume tomorrow. CK

12/26/06, 1100 PST -  Well, we couldn't resist.  We showed up this evening to try a contact even though the day was scheduled for crew rest.  However, extreme winds made it downright scary to be in the station.  We started a contact, but the wind overpowered the dish's AZ servo and made control erratic enough that we terminated track.  Beacon was acquired, but no commanding was attempted.  The whole building is shaking....  CK

12/23-26/06 -  Happy holidays!  The ops team is standing down for some well-earned crew rest over the holiday.  We're planning to be back in action on 12/27.  CK

12/22/06, 0640 PST -  Baseline mission results for GeneSat's biological experiment were distributed to all mission scientists and engineers via the internet approximately 1 hour after the experiment concluded (and less than 30 min after that data had been telemetered to the ground).  In addition, plots of the primary science data were distributed to the public via the internet 90 min after the experiment concluded (and about 1 hour after that data had been telemetered to the ground). CK

 

12/22/06, 0550 PST -  We have successfully downloaded all of the baseline biological data for the 96 hour experiment period.  More data will continue to be collected regarding payload and bus performance in order to meet all mission requirements; however, completion of the baseline biological data set represents a significant milestone for the GeneSat Team. CK

12/22/06, 0400 PST -  GeneSat's biological experiment ends at approximately 0500 PST this morning, and based on operational performance up to this point, the ops team hopes to complete the download of all of the baseline biological results today (we note that the original operational plan prior to launch was to have this done within 4 months after launch; for those of you keeping track, GeneSat has been in orbit less than 1 week). 

Our first contact today was run at 0345 PST; we will be running another one at 0525 PST. In the first contact, we were able to download about 75% of the past day's biological data. During this first contact, only 47 % of transmitted commands were successfully processed and verified.  This is a drop from yesterday's rate of 54%, but still well above the 30% we had been achieving during the first 5 days of operation.  There may have been a weather impact; there is some cloud cover today, and the fog to the north was particularly bad this morning. As previously reported, once the biology data is down and the crew has had a well-deserved holiday rest, we will be conducting a variety of controlled engineering analyses, to include a full characterization of the 2.4 GHz command channel. CK

12/22/06, 0330 PST - NOTE TO HAMS - The 100 hour Grand Prize Competition will conclude at 0900 PST this morning.  There will NOT be any extension to this duration.  To have your received and decoded beacon packets accepted for the competition, they must be submitted via the beacon packet submission web site by 1200 PST on 12/22/06.  It may take a few days to announce the winner, so please stand by.  In the meantime, a posting of the winners of the First Contact Prize and the First 20 Stations Prize is located here. CK

12/21/06, 0800 PST - Two contacts were run this morning at 0430 and 0600 PST.  Bus and payload health are outstanding.  Experiment data for the previous 24 hours of biological growth was retrieved.  The Mission Dashboard includes telemetry snapshots for key bus and payload data points. Use of the 2.4 GHz command channel was dramatically better today than at any previous day of operation - the command channel success rate (fraction of transmitted 
commands that result in successful reception by the vehicle and successful ground reception of the telemetry response) jumped from ~ 30% to slightly above 50%.  We attribute this primarily to the increasing accuracy of the orbit estimate, and has been mentioned before, we hope to see this number continue to rise as the estimate continues to improve.  We note that nearly all commanding has been done in the very early morning hours, so over time we are also interested in assessing how time-varying (and most likely directional) 2.4 GHz noise affects the command link (e.g. how 2.4 GHz wireless noise increases during normal working hours). CK

SCU senior Phelps Williams sets his console on fire with a historic command success rate of 54% during today's GeneSat contacts. [courtey M. Rasay]

12/21/06, 0455 PST - We are happy to report the winner of the GeneSat First Contact Prize for the amateur radio community: Ralph Wallio, WØRPK.  Ralph successfully decoded 5 beacon packets at approximately 05:44:28 PST on 12/16, which was about 85 minutes after GeneSat deployed from the PPOD.  Ralph is located in Iowa.

12/20/06, 0755 PST - Three contacts were run this morning at 0345, 0525, and 0700 PST.  Bus health remains nominal, and biology continues to grow. This is day 2 of the 4 day experiment. Payload temperature control is keeping the biology at the control setpoint.  Beacon comm is outstanding, and more than 5850 beacon packets have been submitted by external HAM operators!  2.4 GHz comm quality continues to fluctuate regarding the conditions for good reception; we continue to have ~30% success rate in successful commanding.  See the Mission Dashboard for the health summary, plots, etc.CK

The mission ops team in the early AM after having successfully turned on the biology experiment over 4 contacts on 12/18/06 [courtesy P. Mahacek (who was too scared to climb the 50 yr old scaffolding...)]

12/19/06, 0800 PST - The mission is proceeding very well.  Three contacts were run today (~0435, 0615 and 0755 PST) in order to download data from the biology experiment as well as to continue monitoring vehicle health.  Biology data shows definite growth over the first of 4 days of expected growth.  Payload temperature control is doing a good job of holding the median payload temperature close to 34 deg C (the latest data shows median payload temp of 33.42). 2.4 GHz communications is functional and supporting the required level of procedural execution by the ops team, but the link continues to be sporadic, with only ~30% success in terms of getting round trip commands and their telemetry responses; we expect improvements in the orbit estimate over time to lead to a better success rate. The beacon downlink (both via SRI and by amateurs throughout the world) is excellent although the frequency has shifted down to about 437.0645 MHz range. CK

12/18/06, 0745 PST - The final contact of the day focused on monitoring the performance of the temperature control system for the experiment.  Nominally, the experiment should run at ~ 34 deg C.  Based on the temperature profile during the experiment start-up sequence (which lasts ~ 96 min), the processor had set the temperature control point to 32 deg C (a somewhat expected situation). The team downloaded payload data to support analysis of the temperature profile during the start-up sequence.  Conditions looked good, and as a result, the set-point was commanded to 34 deg C.  Part of the start-up sequence also involves injecting "feeding" the E. Coli with a glucose solution.  The combination of food and a warm environment should bring them out of stasis such that they will grow over a ~ 96 hour period.  During this growth period, each of the samples will be optically monitored in order to study their growth and their metabolism in the microgravity environment (this will be compared to data from a ground-based experiment in our flight back-up spacecraft).  Bottom line - The beasties have been warmed up and fed!  CK

12/18/06, 0535 PST - The experiment is running! The amateur radio Grand Prize Competition is up and running (packets collected after 0500 PST are valid for the competition) - Good luck to all of you!  CK

12/18/06, 0500 PST - The experiment should now be running.  We will be confirming this during the next contact at 0526 PST. CK

12/18/06, 0430 PST - Amateur radio operators.... The GeneSat experiment is expected to start at 0500 PST (30 min from now).  The Grand Prize Competition will start at that time and will run for approximately 100 hours.  A few caveats.... First - our next contact begins at 0526, and during that contact we will verify whether the experiment has, in fact, been successfully initiated.  If not, we will delay the competition until further notice.  But assuming positive confirmation, the competition is on. Second, the competition is scheduled to run for ~100 hours, which covers the "interesting" part of the biological experiment.  This interesting part, however, is a function of temperature.  So, as the competition progresses, we may extend the competition phase by several hours.  We will post any such extensions here in this log, so stay tuned. Good luck to all, and be sure to submit your packets via the web site to participate.  Details on the competition are here. CK 

12/18/06, 0353 PST - The experiment start command has been successfully scheduled.... the contact is in progress.... CK 

12/18/06, 0347 PST - Our contact is about to start... hey, we have beacon! Gotta go... CK

12/18/06, 0325 PST - We have an official "thumbs up" from the Mission Management team to proceed with our procedure.  If all health and status checks are confirmed and execution of the procedure goes as planned, we'll be on track to have an experiment start at 0500 PST.  Stay tuned.... CK

12/18/06, 0320 PST - The Mission Management Team continues to analyze and discuss bus and payload health data.  So far, the decision is "go" to execute the Start Experiment procedure.  Again, this decision may still be reversed and/or the procedure may not be completed - we will give a definitive "The competition has started" message on this log. CK

12/18/06, 0300 PST - The ops team is currently preparing a command plan that would initiate the experiment at 0500 PST 12/18/06 (1300 GMT). Note that operationally, this consists of sending a "delayed" command during the next contact (which starts ~0348 PST) such that the "start experiment" command executes at 0500 PST.  For those of you interested, the selection of "start" time for the experiment is a function of the "start" process which includes a much tighter (and more power hungry) payload temperature control loop (which begins ~36 min after "start").  Because of power considerations, the "start" time is selected as a function of eclipse times, etc.  

That said, the Mission Management Team is currently discussing, via teleconference, whether to proceed with this plan.  If the decision is "go", it is worth noting that the procedure we must execute to make all of this happen is complex, and given the limited nature of the 2.4 GHz command link (more on that later) we may not be able to execute the entire sequence in time.  So - amateur radio operators - continue to monitor this page.  We will keep you posted as to our plans, and we will let you know when the experiment (and thus the Grand Prize Competition) begins. CK

12/18/06, 0220 PST - The ops team has just completed the first contact of the day in order to assess spacecraft health.  Pending analysis and discussion that will occur with the engineering team over the next 30 min, a determination will be made regarding whether to attempt to initiate the experiment on one of the subsequent contacts.  We will post an update as soon as this decision is made.  Amateur radio operators interested in the Grand Prize Competition (which begins when the experiment is started), stand by..... CK

12/17/06, 1800 PST - Satellite performance and state is positive enough that, pending the health and status of the satellite over the next 8 hours, the GeneSat team may initiate the biological experiment sometime during the early morning (0345-0715 PST) on 12/18/06.  ATTENTION - HAM RADIO CONTEST PARTICIPANTS - the mission's Amateur Radio Grand Prize will be awarded based on beacon packets received during the active biological growth period.  This period begins with the Start Experiment command that we may be sending early morning PST on 12/18/16.  The growth period, and therefore the contest, will nominally last for 96 hours, although that period may be extended depending on the temperature profile during that time.  Updates will be posted here and details of the contest are here.  But get ready!!!  CK

12/17/06, 0400 PST - Our 2.4 GHz command channel works!!!  We're currently using NORAD's new OBJECT C TLEs:

OBJECT C
1 29655U 06058C   06351.19216215  .00100962  00000-0  17267-2  0    26
2 29655 040.0186 073.4702 0004585 153.7796 206.3451 15.50119393    98

This seemed to work like a charm.  Beacon was LOUD and clear immediately - so clear that one of the team walked over to see if our satellite prototype was on by mistake!  Beacon data started flowing as well, and our attempts at 2.4 GHz commanding were immediatley successful!  Details to follow soon. Given good 2.4 GHz comm, the mission is now in Stabilization Phase. CK 

12/16/06, 1315 PST - Beacon data has been coming in throughout the day, with amateurs across the US, in Japan, etc. submitting valid data packets.  We will be making some of the data plots available later today, but so far, the temperature and power continues to be nominal.  NOTE TO HAMS - numerous reports suggest that the beacon frequency seems to have settled in at a frequency several kHz below the launch frequency of 437.075 MHz. These reports put the frequency somewhere in the 437.067 - 437.071 range. CK

12/16/06, 0645 PST - Great news! Initial decode of beacon packets show limited data, but so far that data shows a nominal temperature (ranging from 10-16 degrees C) and power condition! More reports of HAMs hearing the beacon but not well enough to get data... CK

12/16/06, 0615 PST - Based on the launch vector that has been provided, we have a new set of TLEs for all you HAMs out there:

1 99999U Genevec2 06350.62500000 -.00014664  00000-0 -26801-3 0 00000
2 99999 040.0599 076.9488 0017534 172.9505 257.1789 15.48261244000034 

Good luck - and please submit beacon packets to the submission site.  So far, we've had several stations hear the beacon (SRI, SCU, Cal Poly).  Packets have been successfully received by Ralph Wallio, W0RPK. We're decoding and analyzing the packets now.  CK

12/16/06, 0538 PST - GeneSat is alive!  The ops team ran its first contact from SRI from 0528-0537 PST. GeneSat's beacon signal was heard, although no packets were decoded given our lack of knowledge of the orbit.  The team has just now received the launch vector, so we hope to point more precisely on the next pass in about 90 min. The beacon was also heard by the SCU and Cal Poly amateur radio stations. CK

12/16/06, 0420 PST - GeneSat should be deployed from Cal Poly's PPOD!  Go Cal Poly! CK

12/16/06, 0400 PST - LAUNCH!  Everything looks and sounds great!  GeneSat is expected to be deployed at ~0420 PST. CK

12/16/06, 0345 PST - The GeneSat team is assembled at various locations, watching the launch video on the web.  The ops team is at the SRI station, and most of the other team members are at Ames. CK

12/13/06, 1235 PST - It's official - Saturday, 12/16/06, is now the earliest possible launch date.  From that day through 12/22/06, the daily launch window has now been expanded to 0400-0730 PST.

12/13/06, 0900 PST - Saturday, 12/16/06, now appears to be the earliest possible launch date.  Details pending. CK

12/12/06, 0700 PST - Wallops announces that Friday, 12/15/06, is now the earliest possible launch date.  Launch time remains 0400-0700 PST for each day in the current launch window, which has now been extended from 12/15-22/06.  The report on TacSat-2 is that analysis and testing continue of its flight software and its ground-based simulator at Kirtland AFB. Web sites to watch for more information include the NASA TacSat-2 Mission site and the Wallops TacSat-2 Mission site. CK

12/11/06, 1230 PST - Wallops announces that Thursday, 12/14/06, is the earliest possible date for launch. The problem being investigated is being called "a potential inconsistency in the TacSat-2 spacecraft flight software." CK

12/11/06, 0900 PST - Meetings and teleconferences are occurring and will continue throughout the day to determine the severity of the TacSat-2 problem and what the launch options are.  Although there is no official word at this time, a Tuesday 12/12/06 launch seems very doubtful. The GeneSat-1 Mission Team is in the process of assessing their options as a function of the potential delay.  Because GeneSat-1 carries a biological payload, the length of the delay may affect the viability of the biological experiment; several options exist to address this issue. CK

12/10/06, 2245 PST - With only a few hours left in the countdown, the launch has been scrubbed for 12/11/06.  Word is that there is a problem with the primary satellite, TacSat-2.  The current launch period runs from 12/11-20/06 so there are still opportunities to launch in the short term depending on the severity of the problem.  Web sites to watch for more information include the NASA TacSat-2 Mission site and the Wallops TacSat-2 Mission site. CK

12/10/06, 0930 PST - A successful Configuration Control Board (CCB) review has been held to review the upgrade of the SRI Communication Station feed. A quad feed was installed in order to support operation of 4 spacecraft in the next 5 weeks.  CCB approval was given based on positive test data. CK

12/9/06, 1800 PST - This evening's successful Space Shuttle launch clears the way for our launch on Monday morning at 0400 am PST.  The launch can be watched via the web from the Wallops Launch Webcast site.

12/8/06, 1700 PST - The GeneSat Mission Team is using this week to prepare for the launch, currently scheduled for 12/11/06 at 0400 PST.  The Mission Team is actively tracking developments with the Space Shuttle launch.  If the Shuttle does not launch on Saturday night (12/9/06) there is a reasonable chance that it will cause a delay in the GeneSat launch since critical NASA tracking facilities must be used for both missions (and the Shuttle has priority).  Additional SCU students on the operations team are going through a formal training program to become certified to operate the vehicle.  In addition, SCU personnel are now certified to operate the SRI communication station. CK

12/4/06, 1700 PST - Operations Readiness Review (ORR) at NASA Ames Research Center. The team was given Approval To Proceed (ATP) from the ORR review team. CK

 

 

  Santa Clara University Robotics Systems Laboratory

Since 1145 PST 12/20/06 

Hit Counter